Construct a​ 90% confidence​ interval, using the inequality d overbarminust Subscript c Baseline StartFraction s Subscript d Over StartRoot n EndRoot EndFraction less thanmu Subscript dless thand overbarplust Subscript c Baseline StartFraction s Subscript d Over StartRoot n EndRoot EndFraction .

To test the effectiveness of a new drug that is reported to increase the number of hours of sleep patients get during the​ night, researchers randomly select 13 patients and record the number of hours of sleep each gets with and without the new drug.

Patient   Sleep without the drug   Sleep using the drug
1 2.6 3.5
2 4.2   6.0
3 2.2 3.6
4 3.6 4.2
5 4.2 4.9
6 2.4 3.6
7 3.3 4.8
8 3.5 6.1
9 1.9 3.7
10 4.7   5.9
11 5.2 7.1
12 5.3 6.7
13 5.2 7.0

The confidence interval is

nothingless than

Round to two decimal places.
Answer #1
Sample #1 Sample #2 difference , Di =sample1-sample2 (Di - Dbar)²
2.6 3.5 -0.900 0.2983
4.2 6 -1.800 0.1252
2.2 3.6 -1.400 0.0021
3.6 4.2 -0.600 0.7160
4.2 4.9 -0.700 0.5567
2.4 3.6 -1.200 0.0606
3.3 4.8 -1.500 0.0029
3.5 6.1 -2.600 1.3314
1.9 3.7 -1.800 0.1252
4.7 5.9 -1.200 0.0606
5.2 7.1 -1.900 0.2060
5.3 6.7 -1.400 0.0021
5.2 7 -1.800 0.1252
sample 1 sample 2 Di (Di - Dbar)²
sum = 48.3 67.1 -18.8 3.612

mean of difference ,    D̅ =ΣDi / n =   -1.4462
std dev of difference , Sd =    √ [ (Di-Dbar)²/(n-1) =    0.5487
-----------------

sample size ,    n =    13          
Degree of freedom, DF=   n - 1 =    12   and α =    0.1  
t-critical value =    t α/2,df =    1.7823   [excel function: =t.inv.2t(α/2,df) ]      
                  
std dev of difference , Sd =    √ [ (Di-Dbar)²/(n-1) =    0.5487          
                  
std error , SE = Sd / √n =    0.5487   / √   13   =   0.1522
margin of error, E = t*SE =    1.7823   *   0.1522   =   0.2712
                  
mean of difference ,    D̅ =   -1.446          
confidence interval is                   
Interval Lower Limit= D̅ - E =   -1.446   -   0.2712   =   -1.7174
Interval Upper Limit= D̅ + E =   -1.446   +   0.2712   =   -1.1749
                  
so, confidence interval is (   -1.72 < Dbar <   -1.17 )